OpenStudy (anonymous):

A rectangle has an area of 264 square centimeters and a length of 22 centimeters... what is the width

OpenStudy (y1066380):

264/22

OpenStudy (mathmale):

First of all, what is the standard formula for the area of a rectangle of length l and width w? A = ???

OpenStudy (y1066380):

Because the area of rectangle = height*width

OpenStudy (anonymous):

A=b*h

OpenStudy (anonymous):

so I just divide 264 by 22

OpenStudy (mathmale):

Cool: A=bh. We could also write this as A=L*W. This could be solved for W: W=A/L. So: Yes, livetoskate, you need only divide the area 264 by the length 22.

OpenStudy (anonymous):

its 12

OpenStudy (anonymous):

thank you much...

OpenStudy (mathmale):

My pleasure!